Answer the questions below about the quadratic function.
g(x)=2x^2+20x+49
Does the functions have a minimum OR maximum value?
Where does the minimum or maximum value occur?
What IS the function's minimum or maximum value?

Answer by Alan3354(69443)

Answer the questions below about the quadratic function.
g(x)=2x^2+20x+49
Does the functions have a minimum OR maximum value?
It has a min, since it opens upward.
----------------------------
Where does the minimum or maximum value occur?
On the line of symmetry, as always.
-------------
What IS the function's minimum or maximum value?
The line of symmetry is x = -b/2a = -20/4
x = -5
The min is g(-5)
g(-5) = -1
--> (-5,-1)